OpenStudy (anonymous):

A cyclist bikes at a constant speed for 25 miles. He then returns home at the same speed but takes a different route. His return trip takes one hour longer and is 30 miles. Find his speed. 1- 5 mph 2- 7 mph 3- 4 mph 4- 6 mph

OpenStudy (anonymous):

i really want to know how can i do these problem i will give medals for it

OpenStudy (pawanyadav):

Total distance travel=25+30 mile=55 mile

OpenStudy (anonymous):

\[\frac{ 30 }{ x+1}-\frac{ 25 }{ x}=0\]

OpenStudy (anonymous):

and the answer will be 5mph

OpenStudy (anonymous):

\[30x-25x-25=0\]

OpenStudy (pawanyadav):

First convert 55 miles to meter

OpenStudy (anonymous):

\[5x=25\]

OpenStudy (pawanyadav):

And you got x=5 hr

OpenStudy (pawanyadav):

So speed=distance (in meter)/5

OpenStudy (anonymous):

thanks buddy

OpenStudy (pawanyadav):

You are welcome